OpenStudy (anonymous):

if a and b are two irrational numbers , are a+b and a-b rational or irrational?

Directrix (directrix):

Let a = π which is irrational Let b = π which is irrational Is a - b = π - π = ? --> Is the differnce rational or irrational? @shyonaxo